# FGCE: Feasible Group Counterfactual Explanations for Auditing Fairness

### v1

@@ V Experiments
We use the Gender attribute to define groups and denote the created groups $G_{0}$ and $G_{1}$. $G_{0}$ corresponds to females and $G_{1}$ to males. Table VII presents related statistics.
@@ V-A1 Burden
In Table VIII, we report $k_{0}$ and the minimum cost $d_{0}$ for full coverage. In both the Compas and the Adult datasets, $G_{1}$ requires significantly more counterfactuals for full coverage, due to the larger number of WCCs and the difficulty of covering specific WCCs (as indicated in Figure 6). Across all datasets, $G_{0}$ requires less cost for full coverage. This suggests that $G_{0}$ is easier to cover in general as it requires fewer resources in terms of both the number of counterfactuals and the associated cost.
@@
Next, we use the AUC-related fairness measures to account for the various trade-offs between the number $k$ of counterfactuals, cost $d$, and coverage $c$. Table IX shows results of the $kAUC$ plots of coverage against cost, Table IV of the $dAUC$ plots of coverage against the number of counterfactuals, and Table V of the $cAUC$ plots of cost against the number of counterfactuals. The AUC scores in Tables IX and IV are normalized using the maximum values for full coverage, while the AUC scores in Table V are normalized using the $cAUC$ of the plot of the minimum cost value in any of the groups against $k$. For $kAUC$ and $dAUC$, the larger the AUC scores, the better, while for $cAUC$ the smaller the AUC scores, the better. More details about normalization and the selection of the $x$-axis values for all AUC plots are in the supplementary material.
The AUC measures provide a deeper analysis of fairness than $k_{0}$ and $d_{0}$. For the Student dataset, Table IX shows that although $G_{1}$ achieves full coverage with $k=1$, $G_{1}$ requires larger costs than $G_{0}$ to achieve full coverage for $k>1$. For large values of $k$, the $AUC$ scores of both groups become similar. Table IV shows that $G_{1}$ can achieve better coverage for lower costs and smaller $k$ values than $G_{0}$. Finally, if we want to optimize for cost efficiency, Table V shows that $G_{1}$ typically requires fewer resources for coverage up to 75%. For the COMPAS dataset, despite the many WCCs of $G_{1}$, Table IX shows that with just $k=1$, $G_{1}$ achieves over 50% coverage and Table IV that $G_{1}$ achieves similar coverage with comparable costs and comparable $AUC$-scores with $G_{0}$. Table V indicates that $G_{0}$ generally requires fewer counterfactuals and incurs lower costs for all coverage levels but for $c=0.75$. For the Adult dataset, $G_{1}$ gets better $kAUC$-scores than $G_{0}$ only for small values of $k$ (Table IX) and requires a large $k$ for coverage even for large costs (Table IV) indicating that there are factuals in $G_{1}$ that are hard to cover. Table V indicates that $G_{1}$ generally requires many counterfactuals and incurs high costs for all coverage levels.
@@ Appendix C Hyper-parameters
Optimal $k$AUC and $d$AUC Scores The optimal $k$AUC and $d$AUC scores are computed by evaluating the AUC of the values along the x-axis (number of counterfactuals $k$ for $k$AUC, or distances $d$ for $d$AUC) while maintaining the coverage at its maximum. These optimal scores provide a baseline for normalization, ensuring that the $k$AUC and $d$AUC scores reflect the best achievable performance across all $k$ or $d$ values, respectively.
Optimal $c$AUC score The $cAUC$ score is calculated by evaluating $cAUC$ for common $k$ values across groups. The optimal $cAUC$ score is determined by identifying the minimum cost among all groups for each shared $k$, and using this minimum cost for all common $k$ values to compute the metric. This approach ensures that the optimal $cAUC$ reflects the lowest effort required across all groups for a given coverage level. The resulting optimal $cAUC$ score is then used to normalize the $cAUC$ scores.

### v2

@@ V Experiments
We use the Gender attribute to define groups and denote the created groups $G_{0}$ and $G_{1}$. $G_{0}$ corresponds to females and $G_{1}$ to males. Table I presents related statistics.
@@ V-A1 Burden
In Table II, we report $k_{0}$ and the minimum cost $d_{0}$ for full coverage. In both the Compas and the Adult datasets, $G_{1}$ requires significantly more counterfactuals for full coverage, due to the larger number of WCCs and the difficulty of covering specific WCCs (as indicated in Figure 2). Across all datasets, $G_{0}$ requires less cost for full coverage. This suggests that $G_{0}$ is easier to cover in general as it requires fewer resources in terms of both the number of counterfactuals and the associated cost.
@@
Next, we use the $AUC$-related fairness measures to account for the various trade-offs between the number $k$ of counterfactuals, cost $d$, and coverage $c$. Table III shows results of the $kAUC$ plots of coverage against cost, Table IV of the $dAUC$ plots of coverage against the number of counterfactuals, and Table V of the $cAUC$ plots of cost against the number of counterfactuals. The $AUC$ scores in Tables III and IV are normalized using the maximum values for full coverage, while the $AUC$ scores in Table V are normalized using the $cAUC$ of the plot of the minimum cost value in any of the groups against $k$. For $kAUC$ and $dAUC$, the larger the $AUC$ scores, the better, while for $cAUC$ the smaller the $AUC$ scores, the better. More details about normalization and the selection of the $x$-axis values for all $AUC$ plots are in the supplementary material.
The $AUC$ measures provide a deeper analysis of fairness than $k_{0}$ and $d_{0}$. For the Student dataset, Table III shows that although $G_{1}$ achieves full coverage with $k=1$, $G_{1}$ requires larger costs than $G_{0}$ to achieve full coverage for $k>1$. For large values of $k$, the $AUC$ scores of both groups become similar. Table IV shows that $G_{1}$ can achieve better coverage for lower costs and smaller $k$ values than $G_{0}$. Finally, if we want to optimize for cost efficiency, Table V shows that $G_{1}$ typically requires fewer resources for coverage up to 75%. For the COMPAS dataset, despite the many WCCs of $G_{1}$, Table III shows that with just $k=1$, $G_{1}$ achieves over 50% coverage and Table IV that $G_{1}$ achieves similar coverage with comparable costs and comparable $AUC$-scores with $G_{0}$. Table V indicates that $G_{0}$ generally requires fewer counterfactuals and incurs lower costs for all coverage levels but for $c=0.75$. For the Adult dataset, $G_{1}$ gets better $kAUC$-scores than $G_{0}$ only for small values of $k$ (Table III) and requires a large $k$ for coverage even for large costs (Table IV) indicating that there are factuals in $G_{1}$ that are hard to cover. Table V indicates that $G_{1}$ generally requires many counterfactuals and incurs high costs for all coverage levels.
@@ VI Related Work
VII Conclusions
In this paper, we propose FGCE, a novel graph-based framework for group counterfactual generation that addresses limitations in existing methods by incorporating real-world feasibility constraints and managing trade-offs in counterfactual generation. We also introduce novel fairness measures that allow auditing fairness both at the group and the subgroup level, and offering insights on the trade-offs between the cost for achieving recourse, the number of generated counterfactuals, and the coverage of the group. In future work, we plan to further explore the feasibility graph towards providing path-based fairness metrics and to develop appropriate fairness mitigation approaches.
@@ Appendix C Hyper-parameters
Optimal $kAUC$ and $dAUC$ and $cAUC$ scores The optimal $k$AUC and $d$AUC scores are computed by evaluating the AUC of the values along the x-axis (number of counterfactuals $k$ for $k$AUC, or distances $d$ for $d$AUC) while maintaining the coverage at its maximum. These optimal scores provide a baseline for normalization, ensuring that the $k$AUC and $d$AUC scores reflect the best achievable performance across all $k$ or $d$ values, respectively. The $cAUC$ score is calculated by evaluating $cAUC$ for common $k$ values across groups. The optimal $cAUC$ score is determined by identifying the minimum cost among all groups for each shared $k$, and using this minimum cost for all common $k$ values to compute the metric. This approach ensures that the optimal $cAUC$ reflects the lowest effort required across all groups for a given coverage level. The resulting optimal $cAUC$ score is then used to normalize the $cAUC$ scores.